Find the initial temperature of 0.355 kg of glycerin if it is heated by 600 J of heat to produce a final temperature of 91℃.?


1
Expert's answer
2021-06-07T14:13:12-0400

"Q=cm(t_2-t_1)\to t_1=(Q-cmt_2)/cm="


"=(600-2400\cdot0.355\cdot91)/(-2400\cdot0.355)=90.3°" . Answer